1. A highschool teacher at s small private school assigns trigonometry practice problems to be worked via the net. Students must use a password to access the issues and the time of log in and log off is automatically recorded for the teacher. At the end of the week the teacher examines the amount of time each student spent working on the assigned problems. The data provided below in minutes.

15,28,25,48,22,43,49,34,22,33,27,25,

22,20,39. Find the measure of central tendency.


2. Jonny has been playing golf on the weekends for the past three years. Recently, she started keeping track of her recorded scores. Her scores for June and July at her favorite 9-hole (party 36) golf course are provided below.

45,49,42,56,41,36,34,38,41,40,42,41,

39,38,40,39,36,41. Find the measure of central tendency.


1
Expert's answer
2021-12-09T14:18:36-0500

1.


"15,20,22,22,22,25,25,27,"

"28,33,34,39,43,48,49"

"n=15, \sum_ix_i=452"

"mean=\dfrac{1}{n}\sum_ix_i=\dfrac{452}{15}\approx30.1333"

"median=27"

"mode=22 \text{ \ \ appeared 3 times}"

2.


"34, 36, 36, 38, 38, 39, 39, 40, 40,"

"41, 41, 41, 41, 42, 42, 45, 49, 56"

"n=18, \sum_ix_i=738"




"mean=\dfrac{1}{n}\sum_ix_i=\dfrac{738}{18}=41"

"median=\dfrac{40+41}{2}=40.5"

"mode=41 \text{ appeared 4 times}"